OpenStudy (anonymous):

how do i put these graphs in a TI-83 Plus Calculator a. |f(x)| b. f(|x|) c. f(2x) d. 2f(x) e. -f(x) g. f(-x)

OpenStudy (anonymous):

okay well what is f(x)

OpenStudy (anonymous):

1. f(x) =2x+3 2. f(x)=x^2 -5x -3 3.f(x) =2sin(3x) is that what you mean by what is f(x)?

OpenStudy (e.mccormick):

THis is an overall reference to grahing on that calculator: http://www.prenhall.com/divisions/esm/app/graphing/ti83/Graphing/

OpenStudy (anonymous):

Taking your second example \[f(x) = x^2-5x -3\]. Keep in mind that the TI-83 doesn't have the || symbol (at least on the version in my emulator) so you need to use the abs( option under the MATH menu. a. take the absolute value of the whole function \[|f(x)| = abs(x^2-5x-3)\] b. Just take the absolute value of the x's \[f(|x|) = (abs(x))^2-5abs(x)-3\] c. Multiply the x's by 2 \[f(2x)=(2x)^2-5(2x)-3\] d. Multiply the whole function by 2 \[2f(x)=2(x^2-5x-3)\] e. Multiply the whole function by -1 \[-f(x)=-(x^2-5x-3)\] Multiply just the x's by -1 \[f(-x)=(-x)^2-5(-x)-3\]

OpenStudy (anonymous):

i tried putting the first one but it says Error:window range

OpenStudy (anonymous):

what's the range for the graph showing? you probably have to adjust the x and y values

OpenStudy (anonymous):

If you just do a standard -10 to 10 window you should see it

OpenStudy (anonymous):

is it this ? Xmin=1 Xmax= 2.21875 Xscl= 1 Ymin= .49276365 Ymax= 5.49276365 Yscl=1 this is my first time using a graphing calculator so i don't really know how to use it

OpenStudy (anonymous):

your window is too small. Hit the ZOOM button then hit 6 so it will change to the standard window size

OpenStudy (anonymous):

i got it thank you

OpenStudy (anonymous):

ok